# CommonVoice-SpeechRE and RPG-MoGe: Advancing Speech Relation Extraction with a New Dataset and Multi-Order Generative Framework

### v1

@@ Abstract
Index Terms:
Speech Relation Extraction, Multimodal Information Extraction, Cross-modal Alignment, Triple Extraction
@@ I Introduction
Relation Extraction (RE), a fundamental task in information extraction, aims to extract structured knowledge in the form of relational triples (head entity, relation, tail entity) from unstructured data. RE plays a pivotal role in downstream applications such as knowledge graph construction and search engine optimization [1]. Despite its importance, most existing research focuses on TextRE, which extracts relational triples solely from plain text [2, 3, 4].
@@
Overall, SpeechRE is a relatively new research topic and remains underexplored. However, two notable works, LNA-ED [5] and MCAM [6], have already made significant contributions. Wu et al. [5] introduced the SpeechRE task by applying text-to-speech (TTS) to TextRE datasets, creating two synthetic speech benchmarks. They also provided the first SpeechRE baseline, LNA-ED, which uses a CNN-based length adapter to bridge a speech encoder and text decoder. Building on this, Zhang et al. [6] developed two real-human-speech SpeechRE datasets and proposed MCAM, a more powerful model that employs a Multi-Level Cross-Modal Alignment Adapter to align tokens, entities, and sentences across speech and text.
@@
[…]
@@ II-A Speech Relation Extraction
Relation Extraction (RE) is a fundamental task in information extraction that aims to extract structured knowledge in the form of (head entity, relation, tail entity) triplets from unstructured data. Extensive research has established three dominant paradigms for text-based RE (TextRE): (1) sequence labeling approaches [2, 7], (2) table-filling methods [3, 8], and (3) text generation-based frameworks [4].
With the exponential growth of speech data from sources like podcasts and video content, Speech Relation Extraction (SpeechRE) has emerged as a critical yet underexplored task bridging Information Extraction (IE) and Spoken Language Understanding (SLU) [9]. While substantial progress has been made in related tasks like Speech Named Entity Recognition [10, 11, 12], SpeechRE remains nascent. Two seminal works have shaped this field: [5] pioneered the SpeechRE task by synthesizing benchmarks through TTS conversion of TextRE datasets, proposing the LNA-ED model with a CNN-based length adapter to bridge speech encoders and text decoders. Subsequently, [6] advanced the field by constructing the first real-human-speech dataset and developing the MCAM framework featuring hierarchical cross-modal alignment at token, entity, and sentence levels.
@@ II-B Multi-view Prompt Text Generation
Recent work in aspect-based sentiment analysis has shown that leveraging element order diversity in triples [13] or quadruples [14] during training and inference can enhance model performance and generalization. Inspired by this, we are the first to explore the impact of element order diversity in relational triplets on model performance in SpeechRE, a cross-modal text generation task involving both speech and text. This approach distinguishes our work from prior research and opens new avenues for improving SpeechRE through structured data diversity.
@@ III The New Dataset
We present CommonVoice-SpeechRE, a novel dataset derived from the English subset of the Common Voice 17.0 corpus [15]. Common Voice 17.0 is a large-scale, multilingual speech dataset comprising 20,408 validated hours of recordings across 124 languages, contributed by volunteers globally. Released under the CC-0 license, it permits unrestricted use, modification, and redistribution, making it an ideal foundation for secondary annotation tasks such as Speech Relation Extraction (SpeechRE).
Most samples in Common Voice 17.0 are negative examples lacking entities or relations. To identify potential positive samples, we employed a pre-trained BERT NER tagger [^1] https://huggingface.co/flair/ner-english-ontonotes to analyze transcriptions and filter relevant data. We adopted entity and relation type definitions from the ACE04 and ACE05 datasets, crafting a tailored annotation guide. A team of 10 graduate students (all CET-6 certified) manually labeled approximately 20,000 transcriptions using Label Studio [^2] https://labelstud.io. The annotation process involved dividing the data into batches of no more than 1,000 sentences, with 10% randomly selected for verification. Experienced annotators ensured sentence-level accuracy exceeded 95%; otherwise, the batch was re-annotated.
@@ IV-B1 Speech Encoder
Given an input raw speech signal $\boldsymbol{S}$, we first convert it into log-mel spectrogram features $\boldsymbol{X}$. Subsequently, the features $\boldsymbol{X}$ are fed into the Whisper speech encoder [16] to extract high-level speech features $\boldsymbol{H}$ of the speech:
@@ IV-B2 Latet Relation Prediction Head
[…]
@@ IV-B4 Text Decoder
The Text Decoder uses relation prompts and multi-order triplet generation to decode relational triplets. We utilize the pre-trained Whisper decoder [16] for this purpose. The input token sequence to the decoder consists of three parts:
@@ IV-B5 Training and Inference Strategies
[…]
@@
  • Notes: Subscript pipepipe denotes ASR+TextRE pipeline methods; ‘PL-FT’ indicates fine-tuning with pseudo-labeled data; ‘ASR-PTC’ refers to pre-training with ASR data. Subscript oriori represents the original LNA-ED[5]/MCAM[6] backbone: 24-layer Wave2vec encoder + 12-layer BART-large decoder (520M). Subscript whiwhi denotes Whisper [16] backbones: 24-layer encoder/decoder (770M) or 12-layer encoder/decoder (250M).
@@ V-A Datasets & Evaluation Metrics
We conducted experiments on three datasets: CoNLL04-SpeechRE, ReTACRED-SpeechRE and the CommonVoice-SpeechRE dataset proposed in this paper. The CommonVoice-SpeechRE dataset includes diverse real human speech in its training, development, and test sets. For CoNLL04-SpeechRE and ReTACRED-SpeechRE, since the real human speech test set and partial real human speech training set proposed by [6] have not yet been released, we used the fully TTS-generated speech version released by [5]. Detailed statistics of the datasets are provided in Table II. For evaluation metrics, following previous work [5, 6], we used the micro-F1 score to assess the performance of models in entity recognition, relation prediction, and relation triplet extraction. For an entity, relation or triple to be considered correct, it must exactly match its counterpart in the ground truth tags.
@@ V-B Experimental Settings
Our model was implemented using PyTorch-Lightning [^3] https://github.com/Lightning-AI/pytorch-lightning and PyTorch [17], with OpenAI’s Whisper [^4] Whisper has become a standard backbone in speech processing, similar to BERT and BART in NLP. [16] as the backbone, specifically the whisper-small [^5] https://huggingface.co/openai/whisper-small.en (244M) and whisper-medium [^6] https://huggingface.co/openai/whisper-medium.en (769M) versions. We optimized the model parameters using the Adam optimizer with a learning rate of 1e-5, a batch size of 12. Training epochs were set to 50 for CoNLL04-SpeechRE, 20 for ReTACRED-SpeechRE, and 10 for CommonVoice-SpeechRE. For the relation prediction head, we employed a four-layer CNN with 2D convolutions (kernel size = 3) and progressively increasing channel dimensions (16, 32, 64, 128). During inference, the voting threshold $\lambda_{vote}$ for all order views was set to 2. All hyperparameters were tuned on the development set, and the best-performing checkpoint was selected for test set evaluation. Training was conducted on a single NVIDIA A40 GPU, while inference was performed on a single NVIDIA GeForce RTX 4090 GPU.
@@ V-C Baselines
To comprehensively evaluate the performance of our proposed model, we compare it with three categories of competitive baselines: (1) TextRE Models. These models are designed to jointly extract entities and relations from input text. For a fair comparison, following prior works [5, 6], we adopt three strong TextRE models: TP-Linker [3], Spert [2], and REBEL [4] augmented with two recent advances (BiRTE [7] and OD-RTE [8]) Additionally, to explore the potential of large language models (LLMs) in relation extraction, we include GPT-3.5777gpt-3.5-turbo-0125 and GPT-4888gpt-4-turbo-2024-04-09 as baselines, leveraging their in-context learning capabilities for TextRE tasks. (2) Pipeline SpeechRE Models. These models follow a two-stage pipeline: first, an Automatic Speech Recognition (ASR) module transcribes the input speech into text; second, a TextRE module extracts relation triplets from the transcribed text. To ensure a fair comparison, we follow the setup of prior works [5, 6] and employ the pre-trained wav2vec-large model as the ASR module. For the TextRE module, we use the same five TextRE models mentioned above, resulting in seven pipeline models: TP-Linkerpipe, Spertpipe, REBELpipe, BiRTEpipe, OD-RTEpipe, GPT-3.5pipe, and GPT-4pipe. (3) End-to-End SpeechRE Models. These models are designed to directly extract relation triplets from input speech, without the intermediate step of text transcription. Our proposed RPG-MoGe also falls into this category. As baselines, we include two existing end-to-end SpeechRE models: LNA-ED [5] and MCAM [6]. Additionally, to explore the capabilities of recent advancements in speech-based LLMs, we introduce two in-context learning baselines: GPT-4o-audio999gpt-4o-audio-preview-2024-12-17 and Qwen2-audio [^10] Qwen2-Audio-7B-Instruct [18].

### v2

@@ I Introduction
Relation Extraction (RE), a fundamental task in information extraction, aims to extract structured knowledge in the form of relational triples (head entity, relation, tail entity) from unstructured data. RE plays a pivotal role in downstream applications such as knowledge graph construction and search engine optimization [NERandREsurvery]. Despite its importance, most existing research focuses on TextRE, which extracts relational triples solely from plain text [spanbert, tplinker, rebel].
@@
Overall, SpeechRE is a relatively new research topic and remains underexplored. However, two notable works, LNA-ED [LNA-ED] and MCAM [MCAM], have already made significant contributions. Wu et al. [LNA-ED] introduced the SpeechRE task by applying text-to-speech (TTS) to TextRE datasets, creating two synthetic speech benchmarks. They also provided the first SpeechRE baseline, LNA-ED, which uses a CNN-based length adapter to bridge a speech encoder and text decoder. Building on this, Zhang et al. [MCAM] developed two real-human-speech SpeechRE datasets and proposed MCAM, a more powerful model that employs a Multi-Level Cross-Modal Alignment Adapter to align tokens, entities, and sentences across speech and text.
@@
  • Notes: “#Rel”: Number of Relations; “Sam.”: Samples; g: Indicates samples with real-human speech; : Indicates samples with TTS synthetic speech
@@ II-A Speech Relation Extraction
Relation Extraction (RE) is a fundamental task in information extraction that aims to extract structured knowledge in the form of (head entity, relation, tail entity) triplets from unstructured data. Extensive research has established three dominant paradigms for text-based RE (TextRE): (1) sequence labeling approaches [spanbert, BiRTE], (2) table-filling methods [tplinker, odrte], and (3) text generation-based frameworks [rebel].
With the exponential growth of speech data from sources like podcasts and video content, Speech Relation Extraction (SpeechRE) has emerged as a critical yet underexplored task bridging Information Extraction (IE) and Spoken Language Understanding (SLU) [sner_slue]. While substantial progress has been made in related tasks like Speech Named Entity Recognition [sner_en, sner_fr, sner_cn], SpeechRE remains nascent. Two seminal works have shaped this field: [LNA-ED] pioneered the SpeechRE task by synthesizing benchmarks through TTS conversion of TextRE datasets, proposing the LNA-ED model with a CNN-based length adapter to bridge speech encoders and text decoders. Subsequently, [MCAM] advanced the field by constructing the first real-human-speech dataset and developing the MCAM framework featuring hierarchical cross-modal alignment at token, entity, and sentence levels.
@@ II-B Multi-view Prompt Text Generation
Recent work in aspect-based sentiment analysis has shown that leveraging element order diversity in triples [mvp] or quadruples [bvsp] during training and inference can enhance model performance and generalization. Inspired by this, we are the first to explore the impact of element order diversity in relational triplets on model performance in SpeechRE, a cross-modal text generation task involving both speech and text. This approach distinguishes our work from prior research and opens new avenues for improving SpeechRE through structured data diversity.
@@ III The New Dataset
We present CommonVoice-SpeechRE, a novel dataset derived from the English subset of the Common Voice 17.0 corpus [commonvoiceASR]. Common Voice 17.0 is a large-scale, multilingual speech dataset comprising 20,408 validated hours of recordings across 124 languages, contributed by volunteers globally. Released under the CC-0 license, it permits unrestricted use, modification, and redistribution, making it an ideal foundation for secondary annotation tasks such as Speech Relation Extraction (SpeechRE).
Most samples in Common Voice 17.0 are negative examples lacking entities or relations. To identify potential positive samples, we employed a pre-trained BERT NER tagger [^1] https://huggingface.co/flair/ner-english-ontonotes to analyze transcriptions and filter relevant data. We adopted entity and relation type definitions from the ACE04 and ACE05 datasets, crafting a tailored annotation guide. A team of 10 graduate students (all CET-6 certified) manually labeled approximately 20,000 transcriptions using Label Studio [^2] https://labelstud.io [LabelStudio]. The annotation process involved dividing the data into batches of no more than 1,000 sentences, with 10% randomly selected for verification. Experienced annotators ensured sentence-level accuracy exceeded 95%; otherwise, the batch was re-annotated.
@@ IV-B1 Speech Encoder
Given an input raw speech signal $\boldsymbol{S}$, we first convert it into log-mel spectrogram features $\boldsymbol{X}$. Subsequently, the features $\boldsymbol{X}$ are fed into the Whisper speech encoder [whisper] to extract high-level speech features $\boldsymbol{H}$ of the speech:
@@ IV-B2 Latet Relation Prediction Head
CoNLL04-SpeechRE 5 922 231 288 1,283 343 422 11.3s
ReTACRED-SpeechRE 40 33,477 9,350 5,805 58,465 19,584 13,418 12.9s
gCommonVoice-SpeechRE 45 14,557 2,495 2,494 15,948 2,696 2,728 11.6s
  • Notes: : TTS-synthesized speech; g: real human speech. ReTACRED-SpeechRE enumerates all entity pairs as triplets, including “no_relation” type, while the other two datasets only contain positive triplets.
@@ IV-B4 Text Decoder
The Text Decoder uses relation prompts and multi-order triplet generation to decode relational triplets. We utilize the pre-trained Whisper decoder [whisper] for this purpose. The input token sequence to the decoder consists of three parts:
@@ IV-B5 Training and Inference Strategies
CoNLL04-SpeechRE ReTACRED-SpeechRE gCommonVoice-SpeechRE
@@
  • Notes: Subscript pipepipe denotes ASR+TextRE pipeline methods; ‘PL-FT’ indicates fine-tuning with pseudo-labeled data; ‘ASR-PTC’ refers to pre-training with ASR data. Subscript oriori represents the original LNA-ED[LNA-ED]/MCAM[MCAM] backbone: 24-layer Wave2vec encoder + 12-layer BART-large decoder (520M). Subscript whiwhi denotes Whisper [whisper] backbones: 24-layer encoder/decoder (770M) or 12-layer encoder/decoder (250M).
@@ V-A Datasets & Evaluation Metrics
We conducted experiments on three datasets: CoNLL04-SpeechRE, ReTACRED-SpeechRE and the CommonVoice-SpeechRE dataset proposed in this paper. The CommonVoice-SpeechRE dataset includes diverse real human speech in its training, development, and test sets. For CoNLL04-SpeechRE and ReTACRED-SpeechRE, since the real human speech test set and partial real human speech training set proposed by [MCAM] have not yet been released, we used the fully TTS-generated speech version released by [LNA-ED]. Detailed statistics of the datasets are provided in Table II. For evaluation metrics, following previous work [LNA-ED, MCAM], we used the micro-F1 score to assess the performance of models in entity recognition, relation prediction, and relation triplet extraction. For an entity, relation or triple to be considered correct, it must exactly match its counterpart in the ground truth tags.
@@ V-B Experimental Settings
Our model was implemented using PyTorch-Lightning [^3] https://github.com/Lightning-AI/pytorch-lightning and PyTorch [pytorch], with OpenAI’s Whisper [^4] Whisper has become a standard backbone in speech processing, similar to BERT and BART in NLP. [whisper] as the backbone, specifically the whisper-small [^5] https://huggingface.co/openai/whisper-small.en (244M) and whisper-medium [^6] https://huggingface.co/openai/whisper-medium.en (769M) versions. We optimized the model parameters using the Adam optimizer with a learning rate of 1e-5, a batch size of 12. Training epochs were set to 50 for CoNLL04-SpeechRE, 20 for ReTACRED-SpeechRE, and 10 for CommonVoice-SpeechRE. For the relation prediction head, we employed a four-layer CNN with 2D convolutions (kernel size = 3) and progressively increasing channel dimensions (16, 32, 64, 128). During inference, the voting threshold $\lambda_{vote}$ for all order views was set to 2. All hyperparameters were tuned on the development set, and the best-performing checkpoint was selected for test set evaluation. Training was conducted on a single NVIDIA A40 GPU, while inference was performed on a single NVIDIA GeForce RTX 4090 GPU.
@@ V-C Baselines
To comprehensively evaluate the performance of our proposed model, we compare it with three categories of competitive baselines: (1) TextRE Models. These models are designed to jointly extract entities and relations from input text. For a fair comparison, following prior works [LNA-ED, MCAM], we adopt three strong TextRE models: TP-Linker [tplinker], Spert [spanbert], and REBEL [rebel] augmented with two recent advances (BiRTE [BiRTE] and OD-RTE [odrte]) Additionally, to explore the potential of large language models (LLMs) in relation extraction, we include GPT-3.5777gpt-3.5-turbo-0125 and GPT-4888gpt-4-turbo-2024-04-09 as baselines, leveraging their in-context learning capabilities for TextRE tasks. (2) Pipeline SpeechRE Models. These models follow a two-stage pipeline: first, an Automatic Speech Recognition (ASR) module transcribes the input speech into text; second, a TextRE module extracts relation triplets from the transcribed text. To ensure a fair comparison, we follow the setup of prior works [LNA-ED, MCAM] and employ the pre-trained wav2vec-large model as the ASR module. For the TextRE module, we use the same five TextRE models mentioned above, resulting in seven pipeline models: TP-Linkerpipe, Spertpipe, REBELpipe, BiRTEpipe, OD-RTEpipe, GPT-3.5pipe, and GPT-4pipe. (3) End-to-End SpeechRE Models. These models are designed to directly extract relation triplets from input speech, without the intermediate step of text transcription. Our proposed RPG-MoGe also falls into this category. As baselines, we include two existing end-to-end SpeechRE models: LNA-ED [LNA-ED] and MCAM [MCAM]. Additionally, to explore the capabilities of recent advancements in speech-based LLMs, we introduce two in-context learning baselines: GPT-4o-audio999gpt-4o-audio-preview-2024-12-17 and Qwen2-audio [^10] Qwen2-Audio-7B-Instruct [Qwen2-Audio].